Way Cool produces two different models of air conditioners. The company produces the mechanical systems in its components department. The mechanical systems are combined with the housing assembly in its finishing department. The activities, costs, and drivers associated with these two manufacturing processes and the production support process follow. Process Activity Overhead Cost Driver Quantity Components Changeover $ 472,000 Number of batches 830 Machining 304,000 Machine hours 8,700 Setups 231,000 Number of setups 120 $ 1,007,000 Finishing Welding $ 186,000 Welding hours 6,700 Inspecting 221,000 Number of inspections 890 Rework 64,000 Rework orders 120 $ 471,000 Support Purchasing $ 141,000 Purchase orders 537 Providing space 34,000 Number of units 5,190 Providing utilities 69,000 Number of units 5,190 $ 244,000

Additional production information concerning its two product lines follows.

Model 145 Model 212
Units produced 2,100 3,090
Welding hours 2,600 4,100
Batches 415 415
Number of inspections 490 400
Machine hours 3,350 5,350
Setups 60 60
Rework orders 80 40
Purchase orders 358 179


Required:
1. Determine departmental overhead rates and compute the overhead cost per unit for each product line. Base your overhead assignment for the components department on machine hours. Use welding hours to assign overhead costs to the finishing department. Assign costs to the support department based on number of purchase orders.

Components department
Machine hours 0
Number of setups
Finishing department
Welding hours 0
Number of inspections
Support department
Direct materials 0
Rework orders
Model 145 Activity Driver Departmental OH Rate Total Overhead Cost
Components $0
Finishing $0
Support 0
Model 212 Activity Driver Departmental OH Rate Total Overhead Cost
Components $0
Finishing $0
Support 0

Determine the total cost per unit for each product line if the direct labor and direct materials costs per unit are $290 for Model 145 and $190 for Model 212. (Round your intermediate calculations and cost per unit answers to 2 decimal places.)

Model 145 Model 212
Materials and Labor per unit
Overhead cost per unit
Total cost per unit
Answer #1

Solution:

Calculation of Department Overhead Rates:

Components Department on the basis of machine hours

Process Activity

Overhead

Changeover

4,72,000.00

Machining

3,04,000.00

Setups

2,31,000.00

10,07,000.00

Total number of Machine Hours = Machine Hours of Model 145 + Machine Hours of Model 212

= 3,350 + 5,350 = 8,700 Hours

Therefore, Overhead Rate per Hour = $10,07,000 / 8,700 Hours = $115.75 per machine hour

Finishing Department on the basis of welding hours

Process Activity

Overhead

Welding

1,86,000.00

Inspecting

2,21,000.00

Rework

64,000.00

4,71,000.00

Total number of Welding Hours = Welding Hours of Model 145 + Welding Hours of Model 212

= 2,600 + 4,100 = 6,700 Hours

Therefore, Overhead Rate per Hour = $471,000 / 6,700 Hours = $70.30 per welding hour

Support Department on the basis of number of purchase orders

Process Activity

Overhead

Purchasing

1,41,000.00

Providing space

34,000.00

Providing utilities

69,000.00

2,44,000.00

Total number of purchase orders = Purchase orders of Model 145 + Purchase orders of Model 212

= 358 + 179 = 537 purchase orders

Therefore, Overhead Rate per order = $244,000 / 537 orders = $454.38 per purchase order

Calculation of Overhead rate per unit for the product lines.

Model 145

Activity Driver

Department OH Rate

Total Overhead Cost $

Components

3,350.00

115.75

3,87,752.87

Finishing

2,600.00

70.30

1,82,776.12

Support

358.00

454.38

1,62,666.67

Total Overhead Cost $

7,33,195.66

Number of Units produced

2,100

Overhead cost per unit

349.14

Model 212

Activity Driver

Department OH Rate

Total Overhead Cost $

Components

5,350.00

115.75

6,19,247.13

Finishing

4,100.00

70.30

2,88,223.88

Support

179.00

454.38

81,333.33

Total Overhead Cost $

9,88,804.34

Number of Units produced

3,090

Overhead cost per unit

320.00

Calculation of Total Cost per unit

Model 145

Model 212

Materials and Labour per unit $

290.00

190.00

Overhead cost per unit $

349.14

320.00

Total cost per unit $

639.14

510.00